OpenStudy (anonymous):

Solve 6x^(-2)+x^(-1)=2

OpenStudy (anonymous):

Multiply both sides by x^2.

OpenStudy (anonymous):

That is a quadratic equation with solutions -3/2 and 2.

OpenStudy (anonymous):

Shouldn't I multiply by x^-2?

OpenStudy (anonymous):

no multiply by x^2 is correct , giving: 6 + x = 2x^2 2x^2 - x - 6 = 0 u can solve this by factorisation

OpenStudy (anonymous):

\[how \to find the inverse of y = x^3-x+2\]